40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 1 16:06:32 2014, Koji, Configuration, LSC, DRMI locked 6x
    Reply  Sat Nov 8 16:01:32 2014, ericq, Configuration, LSC, DRMI sensing 6x
       Reply  Mon Nov 10 18:23:10 2014, ericq, Configuration, LSC, DRMI sensing REFL11.pngREFL33.pngREFL55.pngREFL165.pngAS55.png
Message ID: 10693     Entry time: Mon Nov 10 18:23:10 2014     In reply to: 10690
Author: ericq 
Type: Configuration 
Category: LSC 
Subject: DRMI sensing 

ARG, I accidentally permuted the digital demod angles. This significantly weakens the argument for believing AS55I is broken... In fact, Jenne and I did some investigations this afternoon that showed that the channel is indeed working. SRX error signal strangeness remains unexplained, however. 

Also, I have yet to compensate for the gain of the violin filters; the actuator calibration numbers I used were for the SUS-LSC FMs, not the LSC FMs where I was injecting. New measurements will be taken soon, as well, since REFL165 is hopefully improved. 

Corrected plots are below. 

REFL11.pngREFL33.png

REFL55.pngREFL165.png

AS55.png

ELOG V3.1.3-